# A universal black-box quantum Monte Carlo approach to quantum phase transitions

### v2

@@ I Introduction
More precisely, we derive QMC estimators for energy susceptibility (ES) [18, 27, 28] and fidelity susceptibility (FS) [4, 9, 17, 18, 19], which can identify quantum critical points in QPTs without relying on order parameters. While related estimators have appeared in other QMC frameworks [18, 17, 19], our derivations and resulting estimators are developed within the relatively recent permutation matrix representation QMC (PMR-QMC), which offers several key advantages over earlier methods [29, 30, 31, 23, 24, 25, 26]. Most relevant to this work, it was recently shown that ergodic PMR-QMC update rules satisfying detailed balance can be generated via a deterministic, automated algorithm for essentially arbitrary Hamiltonians [23, 24, 25]. Furthermore, our estimators are applicable to arbitrarily complex driving terms, as we demonstrate in this work. To the best of our knowledge, neither of these capabilities was supported by previous approaches. Taken together, these features allow our method to be applied directly to a wide range of systems without the need to manually encode laborious, system-specific details.
@@
II Energy and fidelity susceptibilities
@@
that has quantum critical point (QCP) at $\lambda=\lambda_{c},$ denoting the boundary between two quantum phases of a QPT. Henceforth, we refer to $H_{0}$ as the static term and $H_{1}$ as the driving term for convenience. Given the spectral decomposition, $H(\lambda)\ket{\psi_{k}(\lambda)}=E_{k}(\lambda)\ket{\psi_{k}(\lambda)},$ we define the ES [18, 27, 28]
@@
$\chi_{F}(\lambda)=\lim_{\delta\lambda\rightarrow 0}\frac{-2\ln|\langle\psi_{0}(\lambda)|\psi_{0}(\lambda+\epsilon)\rangle|}{\delta\lambda^{2}},$ (3)
@@
which—aside from the pre-factor of $2$—differs from Eq. 5 only by the denominator. By inspection of Eq. 10 and Eq. 8,
@@
With these integral expressions now derived, we can compare the ES and FS, Eqs. 11 and 9 more directly. Evidently, the important difference is that the FS involves a linear factor of $\tau$ in the integrand, which arises from the difference of the square in the denominators of Eqs. 10 and 5. Intuitively, one might expect that leads to the FS having a sharper peak at quantum critical points, and this holds rigorously by performing a standard scaling analysis [28, 47, 44]. Considering $L^{d}$ particles on a $d-$dimensional lattice that undergoes a second-order QPT at critical point $\lambda_{c},$ one can derive the scaling relations [18],
@@
III Numerical demonstrations
Before deriving formal estimators, we demonstrate our method through a series of numerical experiments using open source PMR-QMC code [32] that builds upon prior work [42, 23] (see Appendix A for basics of PMR-QMC). Each model is a spin-$1/2$ system defined in terms of the standard Pauli operators $X$, $Y$, and $Z$. We use the notation $X_{i}$, $Y_{i}$, and $Z_{i}$ to denote Pauli $X$, $Y$, and $Z$ operators acting on spin $i$, respectively.
III.1 A simple two spin model
We begin by studying,
$H(\lambda)=Z_{1}Z_{2}+\gamma(X_{1}+X_{2})+\lambda(Z_{1}+Z_{2}),$ (19)
Figure 1: FS for the two spin model, Eq. (19), considered in Ref. [48] for different inverse temperatures, $\beta$. Data from QMC are points with 95% confidence intervals from bootstrapping over 5 independent random seeds, whereas direct numerical calculations are shown as lines. Naturally, there is no QMC data for $\beta=\infty$.
III.2 A square lattice TFIM
on a square lattice with periodic boundary conditions (PBC). For our purposes, we focus our efforts on a simple $4\times 4$ model at fixed $\lambda=1$ and compute the ES and FS as a function of inverse temperature for either $H_{J}$ or $hH_{h}$ as the “driving term.” Notably, these results are actually not unconditional ES and FS as defined in Eqs. 16 and 17, but rather, the result of restricting computations only to the positive parity subspace of the TFIM. We refer interested readers to Appendix B for details on this methodological trick, but we remark that our approach is a non-trivial convenience of PMR-QMC that highlights its versatility.
Figure 2: We compute ES and FS restricted to the positive parity subspace (see App. B) of a $4\times 4$ square TFIM with PBC and $\lambda=1$ as a function of $\beta$. Our FS results are consistent with a similar computation performed with SSE-QMC [18].
In the $T=0$ limit, this is correct since
follows from plugging in $H_{J}=H-hH_{h}$ into Eqs. 10 and 5 alongside the observation $\langle\psi_{n}|H|\psi_{0}\rangle=0$ for all $n\neq 0$. In these plots, we have chosen $h=1$, and correspondingly, we find that both the ES and FS converge for either perturbation choice ($H_{J}$ or $hH_{h})$ when $\beta$ is large enough. Yet for any finite $\beta$, the relation is less trivial, and in fact, one finds that there is no simple relation that holds across all inverse temperatures [18]. Methodologically, this provides a different way to test for empirical convergence of the finite temperature ES and FS to groundstate values.
So far, we have considered simple models of fixed size. To connect these ideas to the study of QPTs in the infinite system limit, we remark that one can simply compute the ES and FS for increasingly larger system sizes. This is done, for example, in Fig. 4 of Ref. [18]. As expected, this results in an increasingly sharper peak in both the ES and FS as $h$ approaches the known critical point $h_{c}\approx 3.044$ [50]. In addition, one can perform finite size scaling and data collapse in accordance with expected scaling relations, Eqs. 12 and 13, to extract the critical exponent $\nu\approx 0.6301$ [49, 18]. Expressed differently, our method is straightforwardly compatible with studying QPTs in the infinite size limit with enough computational effort, which we show more directly for the XXZ model next.
III.3 A square lattice XXZ model
Figure 3: We compute ES and FS for the XXZ model in Eq. 22 on a square lattice with PBC as a function of temperature, $T$, for different model sizes. Our FS results are consistent with a similar computation performed with SSE-QMC [19].
III.4 A random ensemble
Finally, we consider an ensemble of random 100-spin Hamiltonians generated by random unitary rotations of Eq. (19) (see Appendix C for full details of this family). By unitary invariance, observables for these random models agree with the simple two qubit model despite each $H_{U_{i}}=U_{i}H(\lambda)U_{i}^{\dagger}$ and associated driving term $U_{i}H_{1}U^{\dagger}_{i}$ being quite complicated, containing hundreds of random Pauli terms over 100 spins. Using standard QMC approaches, it is not possible to study this ensemble, but it is straightforward to try in PMR-QMC, and the results agree with known expectations as shown in Fig. 4.
(a) The imaginary time correlator, $G(\tau)$, has the expected symmetry around $\beta/2=10.$
(b) The ES correctly predicts critical points despite the $\beta=20$ curve not quite converging to the $\beta=\infty$ curve.
(c) The FS correctly predicts critical points more sharply than ES, and the $\beta=20$ curve matches the $\beta=\infty$ curve.
Figure 4: Comparison of QMC estimates with exact values for an ensemble of 100–spin Hamiltonians generated by random rotations of Eq. (19). Detailed information on the ensemble can be found in Appendix C. QMC estimates include 95% confidence intervals across 10 models, where each calculation employs a different random seed. In all cases, the QMC estimates show excellent agreement with exact finite $\beta$ curves.
IV Our estimators
@@
Given a Hamiltonian of the form $H(\lambda)=H_{0}+\lambda H_{1}$, we can estimate the ES and FS via Eqs. 16 and 17 provided we can estimate $\langle H_{1}\rangle,\int_{0}^{\beta}\langle H_{1}(\tau)H_{1}\rangle$ and $\int_{0}^{\beta/2}\tau\langle H_{1}(\tau)H_{1}\rangle$. In permutation matrix representation QMC (PMR-QMC) [29, 23, 24, 25], we can derive estimators for each of these quantities independent of the details of $H$, and hence, for arbitrarily complicated driving terms $H_{1}$ [26]. This generality is what allowed us to, using a single estimator code, estimate ES and FS for the variety of models in Section III. In this section, we justify this claim by presenting explicit estimators and discussing their properties and computational complexity.
IV.1 PMR-QMC background and notation
@@
A complete justification for our estimators requires substantial background in the permutation matrix representation (PMR) and PMR-QMC which we discuss in Appendix A (see also Refs. [26, 23]). The main derivations are also technical and lengthy and hence contained in Appendix D. Here, we provide a summary mainly to define the necessary notation and basic ideas.
@@ A PMR-QMC estimator $O_{\mathcal{C}}$ is a function of $\mathcal{C}$ such that
which we denote as $O_{\mathcal{C}}\mathrel{\widehat{=}}\langle O\rangle$ to be read “$O_{\mathcal{C}}$ estimates $\langle O\rangle$” for convenience. Some of the present authors derived and explored the properties of a wide variety of estimators in PMR-QMC generally [26]. The basic idea is that since both $H$ and $O$ can always be decomposed in terms of the same group basis $G$, then one can always derive an expression for $O_{\mathcal{C}}$. In many cases of interest, such as if $O=H_{0}$ or $O=H_{1}$ for any bi-partition of $H$, then this formal expression leads to a faithful estimator in practical simulation [26]. This is why our estimators for the correlator, ES, and FS are independent of the choice of bi-partition.
To discuss our estimators more in-depth, we must introduce some standard PMR-QMC notation. First, we denote |$\ket{z_{k}}\equiv S_{\bm{i}_{k}}\ket{z}$ as the basis state resulting from applying the first $k$ permutations from $S_{\bm{i}_{q}}$. Next, we denote the diagonal energies $E_{z_{k}}\equiv\langle z_{k}|D_{0}|z_{k}\rangle.$ Writing $D_{\bm{i}_{k}}$ as the diagonal associated to $P_{\bm{i}_{k}}$, we can define the off-diagonal hopping term, $D_{(z,S_{\bm{i}_{q}})}\equiv\prod_{k}\langle z_{k}|D_{\bm{i}_{k}}|z_{k}\rangle$. Finally, we denote $e^{-\beta[E_{z_{0}},\ldots,E_{z_{q}}]}$ as the divided difference of the exponential $e^{-\beta x}$ with respect to the multiset of diagonal energies given [56, 57]. These thus defined, we can write the standard PMR-QMC result
@@
IV.2 Pure diagonal estimators
For simplicity, we first assume $H_{1}\propto D_{0}$, i.e., as in the TFIM studied in Section III.2. Using the properties of the PMR-QMC, we can readily derive the estimators,
$E_{z}$ =^$\mathrel{\widehat{=}}\langle D_{0}\rangle$ (26)
@@ IV.2 Pure diagonal estimators
as we show in Appendix D. These estimators involve only sums and products of (1) diagonal energies $E_{z_{j}}$ and (2) divided differences of these diagonal energies—the same quantities that appear in the PMR-QMC weight, Eq. 25. As such, they are both readily computable in a PMR-QMC simulation in $O(1)$ and $O(q^{2})$ time, respectively (see Ref. [26] and Section A.4 to understand the time complexity). Utilizing both of these estimators, we can estimate $G^{D_{0}}(\tau;\lambda;\beta)$ via Eq. (15).
Furthermore, the only non-trivial $\tau$ dependence in these estimators is contained in the divided difference expressions of Eq. 27. Amazingly (see Lemma 1), we can readily derive the exact integral relation
@@
which gives an estimator for $\chi_{E}^{D_{0}}$ (see Eq. (16)) when combined with Eq. 26. As shown in Appendix D, a similar but more complicated derivation yields (see Lemma 3),
@@
Several comments are in order. Firstly, our derivations for ES and FS involve first deriving an estimator for the correlator. We then extend this estimator to the ES and FS by an analytical and exact evaluation of the divided difference integrals, i.e., no numerical integration is required. Secondly, these specific estimators are for $H_{1}\propto D_{0}$, but we can readily derive similar estimators for arbitrary $H_{1}$ (see Section IV.4). Even in the general case, the complexity to estimate the correlator, ES, and FS are $O(q^{2}),O(q^{2})$ and $O(q^{4})$, respectively. Thirdly, the complexity for the ES and FS can be improved in certain special cases as we discuss next.
IV.3 Improved ES and FS estimators
@@
Our proof, given in complete detail in Appendix D, amounts to stating and proving several novel divided difference relations that can be used to simplify Eq. 29 and Eq. 30. While this naively only simplifies the $H_{1}\propto D_{0}$ case, the $H_{1}\propto H-D_{0}$ case follows by linearity and the derivation of a few $O(1)$ estimators for average energy $\langle H\rangle$ and similar quantities.
Stated abstractly, the constraint $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$ may seem restrictive, but recall that this applies to the TFIM and XXZ models studied in Section III. More generally, many well studied models in condensed matter physics have such simple driving terms that are either purely diagonal or purely off-diagonal. As such, we have derived a constant time ES estimator for a wide variety of models.
IV.4 General case estimators
Generally, $H_{1}$ may contain both diagonal and off-diagonal terms, i.e., Eq. 19 or the random ensemble studied in Section III. This is not a fundamental issue for PMR-QMC, as whenever $H$ is in PMR form with $R$ term, $H=\sum_{j}D_{j}P_{j},$ we can always write $H_{1}=\sum_{j}\Lambda_{j}D_{j}P_{j}$ for $\Lambda_{j}$ diagonal matrices by the group structure of the PMR (see Appendix E.1). Observables of this type are said to be in canonical PMR form, and can be estimated correctly [26]. Namely, we can readily estimate $\langle H_{1}\rangle$ and $\langle H_{1}(\tau)H_{1}\rangle$ using expressions that generalize Eqs. 26 and 27.
@@ IV.4 General case estimators
from which we can estimate $\langle H_{1}\rangle$ by linearity. Here, $\delta_{P_{l}}^{(q)}$ equals $1$ if the $q^{\text{th}}$ or final permutation in $S_{\bm{i}_{q}}$ is $P_{l}$ and is 00 otherwise, and we have used the shorthand $\Lambda_{l}(z)\equiv\langle z|\Lambda_{l}|z\rangle.$ As for $\langle H_{1}(\tau)H_{1}\rangle$, we can also derive an estimator for $\langle(\Lambda_{k}D_{k}P_{k})(\tau)\Lambda_{l}D_{l}P_{l}\rangle$ that—up to small technicalities—contains the same $\tau$ dependence as Eq. 27 used to estimate $\langle D_{0}(\tau)D_{0}\rangle.$ Hence, even for arbitrary $H_{1}$, we can extend our estimator for $\langle H_{1}(\tau)H_{1}\rangle$ to estimators for $\chi_{E}^{H_{1}}$ and $\chi_{F}^{H_{1}}$ by our novel divided difference integral relations alone.

### v3

@@ I Introduction
More precisely, we derive QMC estimators for energy susceptibility (ES) [4, 14, 13] and fidelity susceptibility (FS) [22, 60, 52, 4, 57, 55], which can identify quantum critical points in QPTs without relying on order parameters. While related estimators have appeared in other QMC frameworks [4, 52, 57], our derivations and resulting estimators are developed within the relatively recent permutation matrix representation QMC (PMR-QMC), which offers several key advantages over earlier methods [24, 27, 3, 8, 2, 5, 19]. Most relevant to this work, it was recently shown that ergodic PMR-QMC update rules satisfying detailed balance can be generated via a deterministic, automated algorithm for essentially arbitrary Hamiltonians [8, 2, 5]. Furthermore, our estimators are applicable to arbitrarily complex driving terms, as we demonstrate in this work. To the best of our knowledge, neither of these capabilities was supported by previous approaches. Taken together, these features allow our method to be applied directly to a wide range of systems without the need to manually encode laborious, system-specific details.
@@
II Results
Before deriving formal estimators in Section IV, we demonstrate our method through a series of numerical experiments using open source PMR-QMC code [20] that builds upon prior work [45, 8] (see Appendix A for basics of PMR-QMC and Appendix B for basic code instructions). Each model is a spin-$1/2$ system defined in terms of the standard Pauli operators $X$, $Y$, and $Z$.
We use the notation for the ES $\chi_{E}$ and the FS $\chi_{F}$ as defined in “Theoretical background: Energy and fidelity susceptibilities” in Section IV; the PMR-QMC conventions follow “PMR-QMC background and notation” also in Section IV. $X_{i}$, $Y_{i}$, and $Z_{i}$ denote Pauli $X$, $Y$, and $Z$ operators acting on spin $i$, respectively. All error bars and continuous error bands estimate the spread of the plotted quantity over independent QMC runs: Either $2\sigma$ (for thermal averages) or direct 95% percentile intervals (for wall-clock times).
A simple two spin model
Figure 1: FS for the two spin model, Eq. (1), considered in Ref. [67] for different inverse temperatures, $\beta$. Data from QMC are averaged over 5 independent random seeds, whereas direct numerical calculations are shown as lines.
We begin by studying,
$H(\lambda)=Z_{1}Z_{2}+\gamma(X_{1}+X_{2})+\lambda(Z_{1}+Z_{2}),$ (1)
Figure 2: We compute ES and FS restricted to the positive parity subspace (see Appendix C) of a $4\times 4$ square TFIM with PBC and $h=1$ as a function of $\beta$. QMC data are averaged over 100 independent runs. Our FS results are consistent with a similar computation performed with SSE-QMC [4].
A square lattice TFIM
on a square lattice with periodic boundary conditions (PBC). For our purposes, we focus our efforts on a simple $4\times 4$ model at fixed $h=1$ and compute the ES and FS as a function of inverse temperature for either $H_{J}$ or $hH_{h}$ as the “driving term.” Notably, these results are actually not unconditional ES and FS as defined in Eqs. 20 and 21, but rather, the result of restricting computations only to the positive parity subspace of the TFIM. We refer interested readers to Appendix C for details on this methodological trick, but we remark that our approach is a non-trivial convenience of PMR-QMC that highlights its versatility.
Figure 3: We compute ES and FS for the XXZ model in Eq. 4 on a square lattice with PBC as a function of temperature, $T$, for different model sizes. QMC data are averaged over at least 200 independent runs. Our FS results are consistent with a similar computation performed with SSE-QMC [57].
In the $T=0$ limit, this is correct since
follows from plugging in $H_{J}=H-hH_{h}$ into Eqs. 14 and 9 alongside the observation $\langle\psi_{n}|H|\psi_{0}\rangle=0$ for all $n\neq 0$. In these plots, we have chosen $h=1$, and correspondingly, we find that both the ES and FS converge for either perturbation choice ($H_{J}$ or $hH_{h})$ when $\beta$ is large enough. Yet for any finite $\beta$, the relation is less trivial, and in fact, one finds that there is no simple relation that holds across all inverse temperatures [4]. Methodologically, this provides a different way to test for empirical convergence of the finite temperature ES and FS to groundstate values.
So far, we have considered simple models of fixed size. To connect these ideas to the study of QPTs in the infinite system limit, we remark that one can simply compute the ES and FS for increasingly larger system sizes. This is done, for example, in Fig. 4 of Ref. [4]. As expected, this results in an increasingly sharper peak in both the ES and FS as $h$ approaches the known critical point $h_{c}\approx 3.044$ [11]. In addition, one can perform finite size scaling and data collapse in accordance with expected scaling relations, Eqs. 16 and 17, to extract the critical exponent $\nu\approx 0.6301$ [12, 4]. Expressed differently, our method is straightforwardly compatible with studying QPTs in the infinite size limit with enough computational effort, which we show more directly for the XXZ model next.
A square lattice XXZ model
Figure 4: Comparison of QMC estimates with exact values for an ensemble of 100–spin Hamiltonians generated by random rotations of Eq. (1). Detailed information on the ensemble can be found in Appendix D. QMC estimates are averaged across 10 random models. In all cases, the QMC estimates show excellent agreement with exact finite $\beta$ curves. (a) The imaginary time correlator, $G(\tau)$, has the expected symmetry around $\beta/2=10.$ (b) The ES correctly predicts critical points despite the $\beta=20$ curve not quite converging to the $\beta=\infty$ curve. (c) The FS correctly predicts critical points more sharply than ES, and the $\beta=20$ curve matches the $\beta=\infty$ curve.
A random ensemble
Finally, we consider an ensemble of random 100-spin Hamiltonians generated by random unitary rotations of Eq. (1) (see Appendix D for full details of this family). By unitary invariance, observables for these random models agree with the simple two qubit model despite each $H_{U_{i}}=U_{i}H(\lambda)U_{i}^{\dagger}$ and associated driving term $U_{i}H_{1}U^{\dagger}_{i}$ being quite complicated, containing hundreds of random Pauli terms over 100 spins. Using standard QMC approaches, it is not possible to study this ensemble, but it is straightforward to try in PMR-QMC, and the results agree with known expectations as shown in Fig. 4.
Simulation resources and scaling
Figure 5: Simulation time used to generate Fig. 1.
Fig. $N_{\mathrm{p}}$ $\min(T_{\mathrm{wall}})$ $\max(T_{\mathrm{wall}})$
1 5 10s 1m
2 $100$ 15s 14.4hrs
3 200400200-400 3.6hrs 8.8days
4 55 4s 8.2m
Table 1: We summarize the resources used to generate data in this work. Further information is contained in Appendix B.
Figure 6: Estimate of total simulation time scaling for the $L\times L$ XXZ model data in Fig. 3. Timings are given with the following simulation parameters fixed: $N_{\mathrm{therm}}=0$, $N_{\mathrm{meas}}=4\times 10^{8}$, $N_{\mathrm{skip}}=5\times 10^{3}$. Errors bars are estimated over 20 independent runs ($N_{\mathrm{p}}=20)$. We refer readers to Appendix B for more details on the meaning of simulation parameters. (Left) Wall-clock time as a function of $L$ at the critical temperature $T_{c}\approx 0.75$, separated by what is being measured. Simulation time scales roughly as $L^{4}=N^{2}$ in this case. (Right) Wall-clock times to simulate the XXZ model as a function of $T$ for different $L$, with all measurement approaches averaging together.
Numerical demonstrations for Figs. 1 and 4, corresponding to studies of Eq. 1 and its random rotations, were performed on a single 2.3 GHz Intel i9 laptop processor. The remaining experiments involving the TFIM and XXZ models (Figs. 2 and 3) were executed on a University of Southern California high-performance computing cluster. The algorithm is highly memory-efficient [8, 25], with modest requirements even for large-scale simulations (approximately 200 MB of memory per run was sufficient for all calculations reported here). Consequently, our performance analysis focuses primarily on time.
As a QMC method, our approach is naturally suited for lazy parallel execution and exhibits near-perfect strong scaling as the number of cores increases (see, e.g., Fig. 8 of Ref. [8]). Table 1 summarizes the number of MPI threads ($N_{\mathrm{p}}$, approximately corresponding to the number of cores) and wall-clock times used to produce the results in Figs. 1, 2, 3 and 4. Additional details regarding simulation parameters and computational resources are provided in Appendix B.
Next, we briefly discuss the empirical scaling of our algorithm. In Fig. 5, we plot the total simulation time used to generate Fig. 1 in our study of the two qubit model in Eq. 1. Though a simple model, simulation time grows as a function of $\beta$ and near the avoided level crossings, as expected.
In Fig. 6, we repeat our study of the XXZ model at the critical temperature $T_{c}\approx 0.75$ at (a) different system sizes $N=L^{2}$ and (b) with different measurements protocols. We also show the modest wall-clock time scaling as a function of $T$ for different $L$. In our timing experiments, all simulations are run on the same 2.6 GHz AMD EPYC 7513 CPU and simulation parameters are fixed to values listed in the caption, chosen to be in the same range as those used for Fig. 3. The different measurements correspond to either running QMC without performing any measurement, while measuring only the ES, or while measuring only the FS. In all cases, times are similar and the scaling is roughly proportional to $N^{2}=L^{4}$. Given that ES and FS have $O(1)$ and $O(q^{3})$ estimators (see Theorems 1 and 2), this may be surprising, but measurements are performed sparsely (every $N_{\mathrm{skip}}=5000$ MC moves), so wall-clock time is dominated by QMC simulation itself. Further discussion and empirical testing of resource scaling is given in Appendix B.
III Discussion
We hope the results of this work will provide a valuable and versatile tool for condensed matter physicists and researchers studying quantum phase transitions.
IV Methods
We introduce requisite technical background then state and discuss the novel PMR-QMC estimators utilized in our study of QPTs in Section II. Importantly, we state Theorems 1 and 2, which provide estimators for ES and FS applicable to most well-studied condensed matter models. At the cost of some efficiency, we also generalize these theorems to arbitrary Hamiltonians.
Energy and fidelity susceptibilities
@@ IV Methods
[…]
@@
$\chi_{F}(\lambda)=\lim_{\delta\lambda\rightarrow 0}\frac{-2\ln|\langle\psi_{0}(\lambda)|\psi_{0}(\lambda+\delta\lambda)\rangle|}{\delta\lambda^{2}},$ (7)
@@
which—aside from the pre-factor of 22—differs from Eq. 9 only by the denominator. By inspection of Eq. 14 and Eq. 12,
@@
With these integral expressions now derived, we can compare the ES and FS, Eqs. 15 and 13 more directly. Evidently, the important difference is that the FS involves a linear factor of $\tau$ in the integrand, which arises from the difference of the square in the denominators of Eqs. 14 and 9. Intuitively, one might expect that leads to the FS having a sharper peak at quantum critical points, and this holds rigorously by performing a standard scaling analysis [13, 16, 21] and manifests in practical simulations [4, 55]. Considering $L^{d}$ particles on a $d$-dimensional lattice that undergoes a second-order QPT at critical point $\lambda_{c},$ one can derive the scaling relations [4],
@@
Given a Hamiltonian of the form Eq. 5, we can estimate the ES and FS via Eqs. 20 and 21 provided we can estimate $\langle H_{1}\rangle,\int_{0}^{\beta}\langle H_{1}(\tau)H_{1}\rangle$ and $\int_{0}^{\beta/2}\tau\langle H_{1}(\tau)H_{1}\rangle$. In PMR-QMC, we can derive estimators for each of these quantities independent of the details of $H$, and hence, for arbitrarily complicated driving terms $H_{1}$. This generality is what allowed us to, using a single estimator code, estimate ES and FS for the variety of models in Section II. In the following, we justify this claim by presenting explicit estimators and discussing their properties and computational complexity.
PMR-QMC background and notation
@@
A complete justification for our estimators requires substantial background in the permutation matrix representation (PMR) and PMR-QMC which we discuss in Appendix A (see also Refs. [19, 8]). The main derivations are also technical and lengthy and hence contained in Appendix E. Here, we provide a summary mainly to define the necessary notation and basic ideas.
@@ A PMR-QMC estimator $O_{\mathcal{C}}$ is a function of $\mathcal{C}$ such that
which we denote as $O_{\mathcal{C}}\mathrel{\widehat{=}}\langle O\rangle$ to be read $O_{\mathcal{C}}$ estimates $\langle O\rangle$” for convenience. Some of the present authors derived and explored the properties of a wide variety of estimators in PMR-QMC generally [19]. The basic idea is that since both $H$ and $O$ can always be decomposed in terms of the same group basis $G$, then one can always derive an expression for $O_{\mathcal{C}}$. In many cases of interest, such as if $O=H_{0}$ or $O=H_{1}$ for any bi-partition of $H$, then this formal expression leads to a faithful estimator in practical simulation [19]. This is why our estimators for the correlator, ES, and FS are independent of the choice of bi-partition.
[…]
@@
Pure diagonal estimators
For simplicity, we first assume $H_{1}\propto D_{0}$, i.e., as in the TFIM studied in Section II. Using the properties of the PMR-QMC, we can readily derive the estimators,
$E_{z}$ $\mathrel{\widehat{=}}\langle D_{0}\rangle$ (26)
@@
as we show in Appendix E. These estimators involve only sums and products of (1) diagonal energies $E_{z_{j}}$ and (2) divided differences of these diagonal energies—the same quantities that appear in the PMR-QMC weight, Eq. 25. As such, they are both readily computable in a PMR-QMC simulation in $O(1)$ and $O(q^{2})$ time, respectively (see Ref. [19] and Section A.4 to understand the time complexity). Utilizing both of these estimators, we can estimate $G^{D_{0}}(\tau;\lambda;\beta)$ via Eq. (19).
Furthermore, the only non-trivial $\tau$ dependence in these estimators is contained in the divided difference expressions of Eq. 27. Amazingly (see Ref. [66] and Lemma 1 in Section E.3), we can readily derive the exact integral relation
@@
which gives an estimator for $\chi_{E}^{D_{0}}$ (see Eq. (20)) when combined with Eq. 26. As shown in Section E.5, specifically Lemma 3, a similar but more complicated derivation yields,
@@
Several comments are in order. Firstly, our derivations for ES and FS involve first deriving an estimator for the correlator. We then extend this estimator to the ES and FS by an analytical and exact evaluation of the divided difference integrals, i.e., no numerical integration is required. Secondly, these specific estimators are for $H_{1}\propto D_{0}$, but we can readily derive similar estimators for arbitrary $H_{1}$ (see “General case estimators”). Even in the general case, the complexity to estimate the correlator, ES, and FS are $O(q^{2}),O(q^{2})$ and $O(q^{4})$, respectively. Thirdly, the complexity for the ES and FS can be improved in certain special cases as we discuss next.
Improved ES and FS estimators
@@
Our proof, given in complete detail in Appendix E, amounts to stating and proving several novel divided difference relations that can be used to simplify Eq. 29 and Eq. 30. While this naively only simplifies the $H_{1}\propto D_{0}$ case, the $H_{1}\propto H-D_{0}$ case follows by linearity and the derivation of a few $O(1)$ estimators for average energy $\langle H\rangle$ and similar quantities.
Stated abstractly, the constraint $H_{1}\propto D_{0}$ or $H_{1}\propto H-D_{0}$ may seem restrictive, but recall that this applies to the TFIM and XXZ models studied in Section II. More generally, many well studied models in condensed matter physics have such simple driving terms that are either purely diagonal or purely off-diagonal. As such, we have derived a constant time ES estimator for a wide variety of models.
General case estimators
Generally, $H_{1}$ may contain both diagonal and off-diagonal terms, i.e., Eq. 1 or the random ensemble studied in Section II. This is not a fundamental issue for PMR-QMC, as whenever $H$ is in PMR form with $R$ term, $H=\sum_{j}D_{j}P_{j},$ we can always write $H_{1}=\sum_{j}\Lambda_{j}D_{j}P_{j}$ for $\Lambda_{j}$ diagonal matrices by the group structure of the PMR (see Section F.1). Observables of this type are said to be in canonical PMR form, and can be estimated correctly [19]. Namely, we can readily estimate $\langle H_{1}\rangle$ and $\langle H_{1}(\tau)H_{1}\rangle$ using expressions that generalize Eqs. 26 and 27.
@@
from which we can estimate $\langle H_{1}\rangle$ by linearity. Here, $\delta_{P_{l}}^{(q)}$ equals $1$ if the $q^{\text{th}}$ or final permutation in $S_{\bm{i}_{q}}$ is $P_{l}$ and is 0 otherwise, and we have used the shorthand $\Lambda_{l}(z)\equiv\langle z|\Lambda_{l}|z\rangle.$ As for $\langle H_{1}(\tau)H_{1}\rangle$, we can also derive an estimator for $\langle(\Lambda_{k}D_{k}P_{k})(\tau)\Lambda_{l}D_{l}P_{l}\rangle$ that—up to small technicalities—contains the same $\tau$ dependence as Eq. 27 used to estimate $\langle D_{0}(\tau)D_{0}\rangle.$ Hence, even for arbitrary $H_{1}$, we can extend our estimator for $\langle H_{1}(\tau)H_{1}\rangle$ to estimators for $\chi_{E}^{H_{1}}$ and $\chi_{F}^{H_{1}}$ by our novel divided difference integral relations alone.